OpenStudy (vennestudent):

Find the Function Rule (0,1) (1,2) (2,5) (3,10) (4,17) (5,26) I get that it is adding odd numbers in order but how do I write that in an equation form

OpenStudy (zarkon):

it is quadratic

OpenStudy (zarkon):

what is the square of 0,1,2,3,4 and 5

OpenStudy (vennestudent):

I don't know. I'm so confused.

OpenStudy (zarkon):

\[0^2=?\] \[1^2=?\] \[2^2=?\] \[3^2=?\] \[4^2=?\] \[5^2=?\]

OpenStudy (vennestudent):

Oh. Thank you!

OpenStudy (zarkon):

so what is the function?

OpenStudy (vennestudent):

Well I know what those are but how does it relate to the function?

OpenStudy (zarkon):

look at the second coordinate of each order pair and compare it to the squared numbers

OpenStudy (vennestudent):

It's one above the squared numbers 3^2=9 (3,10)

OpenStudy (zarkon):

yes...so can you make a function \(f(x)\) such that \(f(3)=10\)

OpenStudy (vennestudent):

thank you!

OpenStudy (zarkon):

so you know the function?

OpenStudy (vennestudent):

Yes